OpenStudy (anonymous):

16x-10y=10 -8x-6y=6 i have to find x and y by elimination please show work !!

OpenStudy (anonymous):

ok, so first you have to get either the x or the y to cancel each other out, so you multiply the second equation by 2 and get\[-16x-12y=12\] then you combine them, and by doing so, eliminate x. \[-22y=22\] then you solve for y divide by 22 and get \[y=-1\] then you plug y back into the original equation (either one) \[16x - (10\times-1) = 10\] so you get \[16x +10 = 10\] solve for x (subtract 10) \[16x=0\] divide by 16\[x=0\] so y=-1 and x=0

OpenStudy (anonymous):

wow ur good i was wayyyyyyyy off track thanks :)

OpenStudy (anonymous):

no problem is this the method you are used to? i assumed elimination was the same as what we call combination

OpenStudy (anonymous):

yes we are learning all three and elimination is the hardest i think